Find the total charge of a cylinder of Radius R and Length L, carrying a charge density that is proportional to the distance from its axis.


1
Expert's answer
2022-01-26T15:30:51-0500

"\oint\vec Ed\vec A=\frac q{\varepsilon_0},"

"E(2\pi r)L=\frac{\rho \pi r^2L}{\varepsilon_0},"

"E=\frac{\rho r}{2\varepsilon_0}."